The trend line for sales of postcards ( in million) is estimated using annual data from 1980 t0 2004.
The result is st = 10 (1.05)^t.
(a) what is the interpretation of the "10" and the term in parentheses ?
(b) Forecast sales of postcards in 2008.
1
Expert's answer
2018-03-06T09:41:08-0500
(a)
In this given equation, the figure 10 is represented as S0, which is a constant. From this, it can be interpreted that the sale of postcards is at least 10 times from the value of given time period or year.
The term in parentheses is written as:
St = 10 (1.05)^t
In St = In 10 (1.05)^t

(b) Forecast sales of postcards in 2008.
Answer:
For the year 2008, the value of t is 29. So, the estimated sales of postcards in 2008 will be:
S29 = 10 (1.05)^29 = 41.16 million